OpenStudy (anonymous):

Punnett Square Q:

OpenStudy (anonymous):

Use the following punnett Square to answer Questions. Y (yellow), y (green)

OpenStudy (aaronq):

You didn't mention the relationship between the alleles. I'm gonna assume that Y (yellow) is dominant over y (green) - that is YY, Yy give yellow and yy are green. 1. There are 3 yellows, YY, and 2 Yy and only 1 green yy so the probability of picking a yellow seed is \(\dfrac{3}{4}\) 2. From the 3 seeds, only one is YY that will give the same phenotype when crossed with itself (i.e. selfed). So the probability is \(\dfrac{1}{3}\) 3. The only way to not pick a yellow, is if you pick all 3 green. \(P(all~green)=(\frac{1}{4})^3\) \(P(at ~least~1~yellow)=1-P(all~green)=1-(\frac{1}{4})^3=0.984375\) 4. Im not entirely sure for this one, i think because the order doesnt matter, it's y*y*g= \((\dfrac{3}{4})^2*\dfrac{1}{4}=0.140625\)

OpenStudy (anonymous):

@aaronq you're awesome!